OpenStudy (mstv1112):

Why can a subtraction problem be rewritten as an addition problem without changing the outcome? Give a real-world scenario where this might be modeled

OpenStudy (mstv1112):

I came up with one I just need it checked ..... I had 2 dogs they had 3 puppies so now I have 5 dogs I sell 3 and now i'm left with my 2 dogs. 2+3=5-3=2

OpenStudy (mstv1112):

A subtraction problem can be rewritten as an addition problem because subtraction is the opposite of addition

OpenStudy (mstv1112):

@skullpatrol

OpenStudy (anonymous):

5-4=1 5+(-4)=1